# On decomposability and subdifferential of the tensor nuclear norm

### v1

@@ 3.3 Tensor spectral and nuclear norms over subspaces
The orthogonal transformation of tensors discussed in Section 3.2 makes the above statement more intuitive. If $\operatorname{p}_{\mathbb{V}_{k}}$ simply keeps the first several entries unchanged and zeroing the remainder for every $k\in[d]$, to maximize $\langle\mathbf{T},\bigotimes_{k=1}^{d}\mathbf{x}_{k}\rangle$ for a tensor $\mathbf{T}\in\mathbb{T}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)$, an optimal $\mathbf{x}_{k}$ must not have a nonzero entry outside $\mathbb{V}_{k}$ for every $k\in[d]$.
@@
Furthermore, there exists a dual certificate $\mathbf{Z}\in\mathbb{T}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)$ with $\|\mathbf{Z}\|_{\sigma}=1$ such that $\langle\mathbf{T},\mathbf{Z}\rangle=\|\mathbf{T}\|_{*}$.
Proof. We assume that $\mathbf{T}\neq\mathbf{O}$ as otherwise the results hold trivially. This means that $\dim(\mathbb{V}_{k})\geq 1$ for every $k\in[d]$ and further $\mathbb{H}=\bigotimes_{k=1}^{d}(\mathbb{V}_{k}\cap\mathbb{S}^{n_{k}})$ is nonempty. It is easy to see that $\operatorname{conv}(\mathbb{H})$ is convex, compact, and centrally symmetric. Let $\mathbb{H}_{k}\subseteq\mathbb{V}_{k}\cap\mathbb{S}^{n_{k}}$ be any normalized basis of $\mathbb{V}_{k}$ for $k\in[d]$. We have
@@
implying that $\operatorname{sp}\bigl(\operatorname{conv}(\mathbb{H})\bigr)=\mathbb{T}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)$. As a result, $\operatorname{conv}(\mathbb{H})$ must be the unit ball of some norm $\|\bullet\|_{\textnormal{H}}:\mathbb{T}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)\rightarrow\mathbb{R}_{+}$.
@@
Therefore, $\operatorname{p}_{\mathbb{T}((\mathbb{V}_{k})_{k=1}^{d})}(\mathbf{S})\in\mathbb{T}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)$ is also a dual certificate of $\|\mathbf{T}\|_{*}$. $\Box$
@@ 4 Decomposability of the tensor nuclear norm
The decomposability of the matrix nuclear norm has been well understood. It has given rise to many important results in machine learning and statistical estimation [57, Section 10]. Specifically, for two matrices $\mathbf{T},\mathbf{S}\in\mathbb{R}^{n_{1}\times n_{2}}$,
@@
The diagonalization fails to work for higher-order tensors in general. As a result, generalizing the decomposability to tensors remains unclear and unsatisfactory. To the best of our knowledge, the only known result is a weak decomposability [50, Lemma 2.1] that only works for third-order tensors. Specifically, the weak decomposability states that
@@ 4.1 A natural decomposability
includes four out of the eight basic subspaces defined by $(\mathbb{V}_{1},\mathbb{V}_{2},\mathbb{V}_{3})$, each of which spanned by at least two $\mathbb{V}_{k}^{\perp}$’s. While this candidate has only resulted a weak decomposability, it has its own theoretical merits that will be discussed later in Section 5. By making a compromise on the subspace size, it turns out that the full decomposability is indeed possible for some other subspaces. We start with the most restrictive candidate, $\mathbb{T}(\mathbb{V}_{1}^{\perp},\mathbb{V}_{2}^{\perp},\mathbb{V}_{3}^{\perp})$, a natural generalization from the matrix case $\mathbb{T}(\mathbb{V}_{1}^{\perp},\mathbb{V}_{2}^{\perp})$. For comparison, an example of the subspaces of $\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ mentioned earlier is offered in Figure 2.
@@
We remark that the matrix case of Theorem 4.3 has not been seen before, to the best of our knowledge. Both Theorem 4.2 and Theorem 4.3 include and generalize the matrix case on the decomposability of the nuclear norm and spectral norm. It is also worth noting that the decomposability in both Theorem 4.2 and Theorem 4.3 is full, albeit for third-order tensors the subspace $\mathbb{T}^{[3]}\bigl((\mathbb{V}_{k})_{k=1}^{3}\bigr)$ is smaller than $\bigoplus_{|\mathbb{I}|\geq 2,\,\mathbb{I}\subseteq[3]}\mathbb{T}^{\mathbb{I}}\bigl((\mathbb{V}_{k})_{k=1}^{3}\bigr)$ in (8) where only a weak decomposability is possible. That being said, the subspace $\mathbb{T}^{[d]}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)$ is still practically useful in, e.g., analyzing the tensor robust PCA in Section 6.
@@ 4.2 An improved decomposability
The proof of this result will be discussed slightly later.
It is important to remark and in fact easy to observe the monotonicity. The decomposability over $\mathbb{U}_{\mathbb{I}_{1}}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)$ and $\mathbb{U}^{\mathbb{I}_{1}}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)$ implies the decomposability over $\mathbb{U}_{\mathbb{I}_{2}}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)$ and $\mathbb{U}^{\mathbb{I}_{2}}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)$ if $\mathbb{I}_{1}\subseteq\mathbb{I}_{2}$, under which $\mathbb{U}_{\mathbb{I}_{2}}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)\subseteq\mathbb{U}_{\mathbb{I}_{1}}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)$ and $\mathbb{U}^{\mathbb{I}_{2}}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)\subseteq\mathbb{U}^{\mathbb{I}_{1}}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)$. Therefore, Theorem 4.4 is essentially for the strongest case $|\mathbb{I}|=2$ whereas the weakest case is for $\mathbb{I}=[d]$, the result in Theorem 4.2. In a short word, the tensor nuclear norm is decomposable over a pair of subspaces if they have at least two disjoint modes. We say that a mode is disjoint for a pair of subspaces if the projections of the subspaces onto the vector space of that mode are orthogonal to each other.
The above point is particularly clear for the matrix case that is included in Theorem 4.4. Among the four basic subspaces $\mathbb{T}(\mathbb{V}_{1},\mathbb{V}_{2})$, $\mathbb{T}^{\{1\}}(\mathbb{V}_{1},\mathbb{V}_{2})$, $\mathbb{T}^{\{2\}}(\mathbb{V}_{1},\mathbb{V}_{2})$, and $\mathbb{T}^{\{1,2\}}(\mathbb{V}_{1},\mathbb{V}_{2})$, as well as any direct sum of some basic subspaces, only the pair $\mathbb{T}(\mathbb{V}_{1},\mathbb{V}_{2})$ and $\mathbb{T}^{\{1,2\}}(\mathbb{V}_{1},\mathbb{V}_{2})$ and the pair $\mathbb{T}^{\{1\}}(\mathbb{V}_{1},\mathbb{V}_{2})$ and $\mathbb{T}^{\{2\}}(\mathbb{V}_{1},\mathbb{V}_{2})$ have at least two disjoint modes, resulting their full decomposability. If a pair of subspaces share only one disjoint mode, such as $\mathbb{T}(\mathbb{V}_{1},\mathbb{V}_{2})$ and $\mathbb{U}^{\{2\}}(\mathbb{V}_{1},\mathbb{V}_{2})$, nothing can be said about the decomposability; see a simple example below. Therefore, the full decomposability of the tensor nuclear norm over $\mathbb{U}_{\mathbb{I}}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)$ and $\mathbb{U}^{\mathbb{I}}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)$ is in fact in the maximal sense for any $\mathbb{I}\subseteq[d]$ with $|\mathbb{I}|=2$.
@@
We would like to remark that Theorem 4.4 is also important in tensor analysis. As an immediate application, it helps to single out a large class of tensors whose upper bounds of the nuclear norm based on tensor partitions (see [35, Theorem 3.1] and [15, Theorem 3.1]) are tight. Specifically, it is stated in [35, Theorem 3.1] that if a tensor $\mathbf{T}$ is partitioned into any set of subtensors $\{\mathbf{T}_{1},\allowbreak\mathbf{T}_{2},\allowbreak\dots,\allowbreak\mathbf{T}_{m}\}$, then
@@
Therefore, any tensor $\mathbf{T}=\mathbf{T}_{1}+\mathbf{T}_{2}$, where $\mathbf{T}_{1}\in\mathbb{U}_{\mathbb{I}}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)$ and $\mathbf{T}_{2}\in\mathbb{U}^{\mathbb{I}}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)$ with $|\mathbb{I}|=2$, immediately becomes a tight example of the upper bound of (12) for $m=2$. The number of blocks, $m$, can be further increased if we further decompose $\mathbf{T}_{1}$ and/or $\mathbf{T}_{2}$ properly. Besides, this class of tensors already includes all the examples discussed in [35, Section 3.2] as a proper subset. For the same reason, the dual version of Theorem 4.4, i.e., Theorem 4.7 to be presented soon, does a similar job to the tightness of the lower bound of (11).
The proof of Theorem 4.4 follows a similar structure to that of Theorem 4.2. It is an immediate consequence of the following two results, Lemma 4.6 and Theorem 4.7. The former can be shown in the same way as that in the proof of Theorem 4.2, and is thus left to interested readers.

### v2

@@ 3.3 Tensor spectral and nuclear norms over subspaces
The orthogonal transformation of tensors discussed in Section 3.2 makes the above statement more intuitive. If $\operatorname{p}_{\mathbb{V}_{k}}$ simply keeps the first several entries unchanged and zeroing the remainder for every $k\in[d]$, to maximize $\langle\mathbf{T},\bigotimes_{k=1}^{d}\mathbf{x}_{k}\rangle$ for a tensor $\mathbf{T}\in\mathbb{T}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)$, an optimal $\mathbf{x}_{k}$ must not have a nonzero entry outside $\mathbb{V}_{k}$ for every $k\in[d]$. We remark that the Frobenius-norm analogue of the second part of Lemma 3.2, i.e., $\bigl\|\operatorname{p}_{\mathbb{T}((\mathbb{V}_{k})_{k=1}^{d})}(\mathbf{T})\bigr\|_{2}\leq\|\mathbf{T}\|_{2}$, has been established in [65, (3.2)].
@@
Furthermore, there exists a dual certificate $\mathbf{Z}\in\mathbb{T}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)$ with $\|\mathbf{Z}\|_{\sigma}=1$ such that $\langle\mathbf{T},\mathbf{Z}\rangle=\|\mathbf{T}\|_{*}$. Besides, $\bigl\|\operatorname{p}_{\mathbb{T}((\mathbb{V}_{k})_{k=1}^{d})}(\mathbf{T})\bigr\|_{*}\leq\|\mathbf{T}\|_{*}$ for any $\mathbf{T}\in\mathbb{R}^{n_{1}\times n_{2}\times\dots\times n_{d}}$.
Proof. We assume that $\mathbf{T}\neq\mathbf{O}$ as otherwise the results hold trivially. This means that $\dim(\mathbb{V}_{k})\geq 1$ for every $k\in[d]$ and further $\mathbb{H}=\bigotimes_{k=1}^{d}(\mathbb{V}_{k}\cap\mathbb{S}^{n_{k}})$ is nonempty. It is easy to see that $\operatorname{conv}(\mathbb{H})$ is convex, compact, and centrally symmetric with respect to the origin, implying that $\mathbf{O}\in\operatorname{conv}(\mathbb{H})$. Let $\mathbb{H}_{k}\subseteq\mathbb{V}_{k}\cap\mathbb{S}^{n_{k}}$ be any normalized basis of $\mathbb{V}_{k}$ for $k\in[d]$. We have
@@
implying that $\operatorname{sp}\bigl(\operatorname{conv}(\mathbb{H})\bigr)=\mathbb{T}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)$. In fact, $\mathbf{O}$ is an interior point of $\operatorname{conv}(\mathbb{H})$ relative to $\mathbb{T}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)$. If it is not, then the supporting hyperplane theorem (see, e.g., [9, Section 2.5.2]) implies that there exists a supporting hyperplane to $\operatorname{conv}(\mathbb{H})$ at $\mathbf{O}$. This means that $\operatorname{conv}(\mathbb{H})$ is entirely contained in that hyperplane because $\operatorname{conv}(\mathbb{H})$ is centrally symmetric with respect to $\mathbf{O}$, a contradiction to $\operatorname{sp}\bigl(\operatorname{conv}(\mathbb{H})\bigr)=\mathbb{T}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)$. As a result, it follows from [61, Proposition 1.1.8] that $\operatorname{conv}(\mathbb{H})$ must be the unit ball of some norm $\|\bullet\|_{\textnormal{H}}:\mathbb{T}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)\rightarrow\mathbb{R}_{+}$, namely the Minkowski functional [61, Definition 1.1.7] of $\operatorname{conv}(\mathbb{H})$.
@@
Therefore, $\operatorname{p}_{\mathbb{T}((\mathbb{V}_{k})_{k=1}^{d})}(\mathbf{S})\in\mathbb{T}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)$ is also a dual certificate of $\|\mathbf{T}\|_{*}$.
To show the last claim, we observe from the duality and self-adjointness that for any $\mathbf{T}\in\mathbb{R}^{n_{1}\times n_{2}\times\dots\times n_{d}}$,
$\bigl\|\operatorname{p}_{\mathbb{T}((\mathbb{V}_{k})_{k=1}^{d})}(\mathbf{T})\bigr\|_{*}$ $=\max_{\|\mathbf{Z}\|_{\sigma}\leq 1}\langle\operatorname{p}_{\mathbb{T}((\mathbb{V}_{k})_{k=1}^{d})}(\mathbf{T}),\mathbf{Z}\rangle$
$=\max_{\|\mathbf{Z}\|_{\sigma}\leq 1}\langle\mathbf{T},\operatorname{p}_{\mathbb{T}((\mathbb{V}_{k})_{k=1}^{d})}(\mathbf{Z})\rangle$
$\leq\max_{\|\operatorname{p}_{\mathbb{T}((\mathbb{V}_{k})_{k=1}^{d})}(\mathbf{Z})\|_{\sigma}\leq 1}\langle\mathbf{T},\operatorname{p}_{\mathbb{T}((\mathbb{V}_{k})_{k=1}^{d})}(\mathbf{Z})\rangle$
$\leq\max_{\|\mathbf{Y}\|_{\sigma}\leq 1}\langle\mathbf{T},\mathbf{Y}\rangle$
$=\|\mathbf{T}\|_{*},$
where the first inequality is due to $\bigl\|\operatorname{p}_{\mathbb{T}((\mathbb{V}_{k})_{k=1}^{d})}(\mathbf{Z})\bigr\|_{\sigma}\leq\|\mathbf{Z}\|_{\sigma}\leq 1$ by Lemma 3.2. $\Box$
@@ 4 Decomposability of the tensor nuclear norm
The decomposability of the matrix nuclear norm has been well understood. It has given rise to many important results in machine learning and statistical estimation [67, Section 10]. Specifically, for two matrices $\mathbf{T},\mathbf{S}\in\mathbb{R}^{n_{1}\times n_{2}}$,
@@
The diagonalization fails to work for higher-order tensors in general. As a result, generalizing the decomposability to tensors remains unclear and unsatisfactory. To the best of our knowledge, the only known result is a weak decomposability [55, Lemma 2.1] that only works for third-order tensors. Specifically, the weak decomposability states that
@@ 4.1 A natural decomposability
includes four out of the eight basic subspaces defined by $(\mathbb{V}_{1},\mathbb{V}_{2},\mathbb{V}_{3})$, each of which is spanned by at least two $\mathbb{V}_{k}^{\perp}$’s. While this candidate has only resulted in a weak decomposability, it has its own theoretical merits that will be discussed later in Section 5. By making a compromise on the subspace size, it turns out that the full decomposability is indeed possible for some other subspaces. We start with the most restrictive candidate, $\mathbb{T}(\mathbb{V}_{1}^{\perp},\mathbb{V}_{2}^{\perp},\mathbb{V}_{3}^{\perp})$, a natural generalization from the matrix case $\mathbb{T}(\mathbb{V}_{1}^{\perp},\mathbb{V}_{2}^{\perp})$. For comparison, an example of the subspaces of $\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ mentioned earlier is offered in Figure 2.
@@
We remark that the matrix case of Theorem 4.3 has been established in [64, (2.2.10)] and as a special case of [38, (1.2)]. Both Theorem 4.2 and Theorem 4.3 include and generalize the matrix case on the decomposability of the nuclear norm and spectral norm. It is also worth noting that the decomposability in both Theorem 4.2 and Theorem 4.3 is full, albeit for third-order tensors the subspace $\mathbb{T}^{[3]}\bigl((\mathbb{V}_{k})_{k=1}^{3}\bigr)$ is smaller than $\bigoplus_{|\mathbb{I}|\geq 2,\,\mathbb{I}\subseteq[3]}\mathbb{T}^{\mathbb{I}}\bigl((\mathbb{V}_{k})_{k=1}^{3}\bigr)$ in (9) where only a weak decomposability is possible. That being said, the subspace $\mathbb{T}^{[d]}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)$ is still practically useful in, e.g., analyzing the tensor robust PCA in Section 6.
@@ 4.2 An improved decomposability
The proof of Theorem 4.4 follows a similar structure to that of Theorem 4.2 and is an immediate consequence of Lemma 4.6 and Theorem 4.7 to be introduced soon. We shall include it after the proof of Theorem 4.7.
It is important to remark and in fact easy to observe the monotonicity. The decomposability over $\mathbb{U}_{\mathbb{I}_{1}}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)$ and $\mathbb{U}^{\mathbb{I}_{1}}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)$ implies the decomposability over $\mathbb{U}_{\mathbb{I}_{2}}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)$ and $\mathbb{U}^{\mathbb{I}_{2}}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)$ if $\mathbb{I}_{1}\subseteq\mathbb{I}_{2}$, under which $\mathbb{U}_{\mathbb{I}_{2}}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)\subseteq\mathbb{U}_{\mathbb{I}_{1}}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)$ and $\mathbb{U}^{\mathbb{I}_{2}}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)\subseteq\mathbb{U}^{\mathbb{I}_{1}}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)$. Therefore, Theorem 4.4 is essentially for the strongest case $|\mathbb{I}|=2$ whereas the weakest case is for $\mathbb{I}=[d]$, the result in Theorem 4.2. In short, the tensor nuclear norm is decomposable over a pair of subspaces if they have at least two disjoint modes. We say that a mode is disjoint for a pair of subspaces if the projections of the subspaces onto the vector space of that mode are orthogonal to each other.
The above point is particularly clear for the matrix case that is included in Theorem 4.4. Among the four basic subspaces $\mathbb{T}(\mathbb{V}_{1},\mathbb{V}_{2})$, $\mathbb{T}^{\{1\}}(\mathbb{V}_{1},\mathbb{V}_{2})$, $\mathbb{T}^{\{2\}}(\mathbb{V}_{1},\mathbb{V}_{2})$, and $\mathbb{T}^{\{1,2\}}(\mathbb{V}_{1},\mathbb{V}_{2})$, as well as any direct sum of some basic subspaces, only the pair $\mathbb{T}(\mathbb{V}_{1},\mathbb{V}_{2})$ and $\mathbb{T}^{\{1,2\}}(\mathbb{V}_{1},\mathbb{V}_{2})$ and the pair $\mathbb{T}^{\{1\}}(\mathbb{V}_{1},\mathbb{V}_{2})$ and $\mathbb{T}^{\{2\}}(\mathbb{V}_{1},\mathbb{V}_{2})$ have at least two disjoint modes, resulting in their full decomposability. If a pair of subspaces share only one disjoint mode, such as $\mathbb{T}(\mathbb{V}_{1},\mathbb{V}_{2})$ and $\mathbb{U}^{\{2\}}(\mathbb{V}_{1},\mathbb{V}_{2})$, nothing can be said about the decomposability; see a simple example below. Therefore, the full decomposability of the tensor nuclear norm over $\mathbb{U}_{\mathbb{I}}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)$ and $\mathbb{U}^{\mathbb{I}}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)$ is in fact in the maximal sense for any $\mathbb{I}\subseteq[d]$ with $|\mathbb{I}|=2$.
@@
We would like to remark that Theorem 4.4 is also important in tensor analysis. As an immediate application, it helps to single out a large class of tensors whose upper bounds of the nuclear norm based on tensor partitions (see [40, Theorem 3.1] and [16, Theorem 3.1]) are tight. Specifically, it is stated in [40, Theorem 3.1] that if a tensor $\mathbf{T}$ is partitioned into any set of subtensors $\{\mathbf{T}_{1},\allowbreak\mathbf{T}_{2},\allowbreak\dots,\allowbreak\mathbf{T}_{m}\}$, then
@@
Therefore, any tensor $\mathbf{T}=\mathbf{T}_{1}+\mathbf{T}_{2}$, where $\mathbf{T}_{1}\in\mathbb{U}_{\mathbb{I}}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)$ and $\mathbf{T}_{2}\in\mathbb{U}^{\mathbb{I}}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)$ with $|\mathbb{I}|=2$, immediately becomes a tight example of the upper bound of (13) for $m=2$. The number of blocks, $m$, can be further increased if we further decompose $\mathbf{T}_{1}$ and/or $\mathbf{T}_{2}$ properly. Besides, this class of tensors already includes all the examples discussed in [40, Section 3.2] as a proper subset. For the same reason, the dual version of Theorem 4.4, i.e., Theorem 4.7 to be presented soon, does a similar job to the tightness of the lower bound of (12).
We are now in a position to state Lemma 4.6, which can be shown in the same way as that in the proof of Theorem 4.2, and is thus left to interested readers.
@@
We are now ready to prove Theorem 4.4.
Proof of Theorem 4.4. The condition of Theorem 4.4 is the same to the condition of Theorem 4.7 and implies the condition of Lemma 4.6. Therefore, by Theorem 4.7, the spectral norm is decomposable over $\mathbb{U}_{\mathbb{I}}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)$ and $\mathbb{U}^{\mathbb{I}}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)$, which implies that the nuclear norm is decomposable over $\mathbb{U}_{\mathbb{I}}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)$ and $\mathbb{U}^{\mathbb{I}}\bigl((\mathbb{V}_{k})_{k=1}^{d}\bigr)$ by Lemma 4.6. $\Box$